# Composable Finite-Size Security of High-Dimensional Quantum Key Distribution Protocols

### v1

@@ VI.2 Variable-length key rates
Next, we come to the variable-length scenario. To illustrate the advantage of finite-size key rates, we need to employ a noise model that translates certain physical parameters of the used devices, such as the pair-production rate of the source or the dark-count rate and the efficiency of the detector and channel parameters like loss and the number of environmental photons (for example solar photons) to coincidence click matrices. The latter can be used to derive the observations for our witness operators (observables). For illustration purposes, we use the noise model developed in Ref. [27], although, of course, the arguments work for any noise model. Additionally, one needs to simulate natural channel fluctuations, as they appear under any realistic conditions. In Appendix C, we discuss and model two different types of channels that may appear in real experiments. We call the first one ‘Statistically Fluctuating Channel’, which refers to a well-characterized, stable channel with channel parameters probability of photon loss $p_{L}$ and a number of excess (solar) photons $n_{\mathrm{sol}}$ that follow a normal distribution around certain characterized means. This represents a stable atmospheric channel under constant weather conditions. The second channel is called the ’Rapidly Fluctuating Channel‘ and describes an atmospheric channel under unstable weather conditions, such as cloudy or foggy nights, or changing brightness, e.g., due to the rising sun. We apply a simplified model and assume that randomly one out of nine Statistically Fluctuating Channels with different means $\mu_{\mathrm{sol}}$ and $\mu_{\mathrm{loss}}$ manifests. This, of course, is only for demonstration purposes, while in reality a plethora of conditions may apply. While the second scenario is more realistic, already the first one suffices to demonstrate the clear advantage of variable-length protocols. Throughout the whole chapter, we fix the dimension to $d=16$. We direct the interested reader to Appendix B for details about our choices for the security parameters. In this chapter, we set $r=\frac{1}{1000}=s$, which turned out to be a sufficiently good choice.
@@
We start our discussion by reviewing the notion of acceptance testing for fixed-length protocols. There, before the protocol is executed, one has to determine and fix an expected channel behavior by determining expected observations for the witness operators. This usually is a single data point. In this case, of course, every tiny deviation from the expected behavior leads to a deviation from the observed statistics. During the acceptance test, one compares these observed statistics with the expected statistics and aborts in case they do not match. Therefore, under this ‘unique-acceptance’ scenario, one would abort the protocol almost all the time, making it secure but essentially useless. We can fix this problem by introducing some margin to extend the set of accepted statistics (see parameter $t_{X}$ in Theorem 1). As elaborated on earlier, by doing so, we build a multi-dimensional interval centered on our original expected statistics, which increases the probability of acceptance. This, however, comes at the cost of extending the set of states that pass the acceptance set, hence the set of states that potentially produced the observed statistics, and thus lowers the key rate. To ease the illustration, we parametrize $t_{X}$ for each observable $X\in\Theta$ by a positive real number $t_{F}$, $t_{X}=t_{F}\mu_{X}$. In contrast, the variable-length security argument is built around the observed statistics and not around the expected statistics and adapts the secure key length to the quality of the observations. Therefore, for variable-length key rates, we can set $t_{F}=0$. As derived in Section IV, the acceptance testing security parameter $\epsilon_{\mathrm{AT}}$ differs between the fixed-length and the variable-length scenario, where $\epsilon_{\mathrm{AT}}$ is the sum of the acceptance-testing epsilons $\epsilon_{X}$ for each single observable $X\in\Theta$. To simplify application for the variable-length scenario, from now on, we require that all the $\epsilon_{X}$ are equal, hence $\epsilon_{\mathrm{AT}}=|\Theta|\epsilon_{X}$ and $\mu_{X}=\sqrt{\frac{x^{2}}{2m_{X}}\ln\left(\frac{2}{\epsilon_{X}}\right)}$.
@@
Let us begin by developing an intuition for those two different notions of key rates on fluctuating channels. In Figure 3a, we simulate $100$ runs of the ‘Statistically Fluctuating Channel’ with the following channel parameters $\mu_{\mathrm{sol}}=2\times 10^{4}$, $\sigma_{\mathrm{sol}}=1\times 10^{4}$, $\mu_{\mathrm{loss}}=0.99$ and $\sigma_{\mathrm{loss}}=0.005$. The pair production rate was set to $0.1$ per time-frame, the detection efficiency was assumed to be $\eta_{D}=0.90$ and the dark-count rate was assumed to be $100$ per second. Furthermore, we set $t_{F}=0.5$. We plot the secure key rate for each run both for the fixed-length as well as for the variable-length protocol. While the fixed-length protocol either produces a key of fixed length or aborts, it can be seen readily how the key lengths for the variable-length protocol adapt to the observed data, while it only fails for two out of $100$ data points, yielding zero key. In the case of acceptance, the difference in key rate between the fixed-length and the variable-length case is (in this example) about $20\%$. However, since for the fixed-length scenario more than half of the runs are not accepted, this results in the average variable-length key rates being roughly $2.5$ higher than the fixed-length key rates.
@@
In real scenarios, the channel parameters are usually unknown, so, for the present example, we chose $t_{F}$ heuristically. This leaves the question of whether in case the behavior of the channel, including its statistical parameters, would be known, one could find the optimal $t_{F}$ for the fixed-length scenario and close the gap in performance. We are going to explore this question in what follows.
@@
In Figure 3b, we examine the key rates for the fixed-length protocol and its acceptance ratio for different values of $t_{F}$. Again, we assumed a ‘Statistically Fluctuating Channel’ with the channel parameters mentioned above and simulated for each datapoint $1000$ runs. The blue dashed line represents the ’one-shot‘ fixed-length key rates, which are the fixed-length key rates upon acceptance, while the light-red dotted curve shows the associated ratio of accepted rounds. The dot-dashed red curve represents the average fixed-length key rate for each data point (which is the product of the one-shot value with the acceptance ratio). For comparison reasons, we also plot the (average) variable-length key rate (black dotted line). Note that the variable-length key rate does not require $t_{F}>0$ and thus is constant with $t_{F}$. One can see that in this example, the one-shot fixed-length key rates are higher than the variable-length key rates for $t_{F}<0.25$. However, due to the low acceptance ratio, the expected fixed-length key rates always remain significantly lower than the variable-length rates. For higher values of $t_{F}$, the acceptance rate slowly converges towards $1$ while at the same time, the one-shot key rate drops, resulting overall in a decreasing average key rate. The highest fixed-length key rate is observed for $t_{F}=1$. However, the exact location of the maximum depends on the channel parameters. Thus, even in case one manages to determine the optimal $t_{F}$ a-priori, the obtained average fixed-length key rates would still be significantly below the variable-length rates.
@@
Already in the case of stable channels with parameters fluctuating statistically around fixed mean values, variable length key rates outperform their fixed-length equivalents in every aspect. In case, in addition, the mean values fluctuate rapidly, the fixed-length acceptance test will only pass in a tiny fraction of all runs, leading to an expected key rate of nearly zero, while the variable-length key rates can to the observed statistics and will produce non-zero key in many of those cases. We illustrate this in Figure 4, where we simulate the chaotic channel with parameters $(p_{\textrm{loss}},n_{\text{sol}})\in\{0.97,0.98,0.99\}\times\{1\times 10^{4},2\times 10^{4},1\times 10^{5}\}$ (see Appendix C for details), where each channel parameter is chosen randomly according to uniform distribution and $t_{F}=0.5$. As expected, the fixed-length protocol aborts in the majority of all cases, yielding zero key. In contrast, the variable-length protocol produces non-negative key for most runs. The average variable-length rates are now about $6.7$ times higher than the average fixed-length key rate. This illustrates that, in real free-space channels, the advantage of the variable-length security argument will be even more striking.
@@ VII Discussion
In recent years, the development of implementations for high-dimensional QKD on the one side and security analyses on the other side proceeded at a vastly different pace. While highly entangled states have been produced in various degrees of freedom [18, 19, 20, 51, 52, 53, 33, 54, 34, 35, 36], theoretical analyses were lagging behind, either relying on practically infeasible measurements [24, 22] or using computationally intense convex optimization methods [27] that are limited to very low dimensions, despite taking hours to days for computation. While this problem was solved in the asymptotic regime by Ref. [31], the practically relevant finite-size regime for HD-QKD was still missing.
In this work, we closed this gap, proving the first composable finite-size security argument for a general high-dimensional QKD protocol against both collective and coherent attacks. The obtained rates converge to the corresponding asymptotic rates [31] in the limit of large $N$. We obtain non-negative key rates already for practically achievable block sizes of $N=10^{8}$ under viable conditions. This shows that HD-QKD is suitable for loss- and noise regimes [55, 56] present in free-space and LEO satellite QKD links.
In addition to the common fixed-length key rates, we also provide practically more useful variable-length key rates, which build the composability argument around the observed statistics rather than the expected statistics, which improves both the practicality and the key rates of the protocol significantly. We discuss and compare those two different notions of security with each other to illustrate the advantages of variable-length key rates.
While we used a time-bin implementation for demonstration purposes, our work is general and can be applied to a variety of platforms with little to no adaptations. In summary, this work provides the missing tool for the full security analysis of general high-dimensional QKD protocols and unlocks the practical use of high-dimensionally entangled states in Quantum Key Distribution. The obvious next steps consist of the application to real implementations and further improvements in the key rate.
@@ Acknowledgements.
F.K. thanks Devashish Tupkary for inspiring discussions and precious feedback on an earlier version of this manuscript, Ian George for insightful discussions and Matej Pivoluska for valuable comments on an earlier version of the manuscript. This work has received funding from the Horizon-Europe research and innovation programme under grant agreement No 101070168 (HyperSpace).
@@ Appendix A Construction of the set $V(\vec{F}^{\mathrm{obs}})$
In the fixed-length scenario, the acceptance test was built upon defining expected statistics and finding the set of density matrices that could have produced these statistics only with some small probability; the complement of this set was the set $\mathcal{S}_{\mathrm{AT}}$. In the variable-length scenario, the construction of the acceptance set follows a slightly different logic. A fixed but otherwise unknown quantum state gives rise to observed statistics $\vec{F}^{\mathrm{obs}}$, which is a random variable. Based on this observation, we aim to construct a set $V(\vec{F}^{\mathrm{obs}})$ such that $\mathrm{Pr}\left[\tau_{AB}\in V(\vec{F}^{\mathrm{obs}})\right]\geq 1-\epsilon_{\mathrm{AT}}$. We are using observables, thus we have to slightly adapt the argument from [39]. Let $\Theta$ denote the set of observables used in the protocol and let $x:=||X||_{\infty}=||X||_{\infty}||\rho||_{1}\geq||X\rho||_{1}$ for $X\in\Theta$. For each $X\in\Theta$, Hoeffding’s inequality [46] tells us
@@ Appendix C Simulation of the Atmospheric Channel
In order to demonstrate the advantages of a variable-length security argument, we simulate secure key rates for two different channel behaviors. First, we consider a statistically fluctuating channel, followed by a rapidly fluctuating channel. Both simulations use the noise model from [27] to relate the experimental quantities pair-production rate $c_{\mathrm{Prod}}$, dark-count rate $d_{\mathrm{dark}}$, detection efficiency $\eta_{d}$, channel loss probability $p_{L}$ and number of environmental (solar) photons $n_{\mathrm{sol}}$ to click matrices. For details regarding this noise model, we refer interested readers to Ref. [27].

### v2

@@ VI.2 Variable-length key rates
Next, we come to the variable-length scenario. To illustrate the advantage of finite-size key rates, we need to employ a noise model that translates certain physical parameters of the devices used, such as the pair-production rate of the source, the dark-count rate, and efficiency of the detector and channel parameters like loss and the number of environmental photons (for example solar photons) into coincidence click matrices. The latter can be used to derive the observations for our witness operators (observables). For illustration purposes, we use the noise model developed in Ref. [BKPH_2023], although, of course, the arguments apply to any noise model. Additionally, one needs to simulate natural channel fluctuations, as they appear under any realistic conditions. In Appendix C, we discuss and model two different types of channels that may appear in real experiments. We call the first one ‘Statistically Fluctuating Channel’, which refers to a well-characterized, stable channel with channel parameters probability of photon loss $p_{L}$ and a number of excess (solar) photons $n_{\mathrm{sol}}$ that follow a normal distribution around certain characterized means. This represents a stable atmospheric channel under constant weather conditions. The second channel is called the ’Rapidly Fluctuating Channel‘ and describes an atmospheric channel under unstable weather conditions, such as cloudy or foggy nights, or changing brightness, e.g., due to the rising sun. We apply a simplified model and assume that randomly one out of nine Statistically Fluctuating Channels with different means $\mu_{\mathrm{sol}}$ and $\mu_{\mathrm{loss}}$ manifests. This, of course, is only for demonstration purposes, while in reality, a plethora of conditions may apply. While the second scenario is more realistic, the first one already suffices to demonstrate the clear advantage of variable-length protocols. Throughout the whole chapter, we fix the dimension to $d=16$. We direct the interested reader to Appendix B for details about our choices for the security parameters. In this section, we set $r=\frac{1}{1000}=s$, which turns out to be a sufficiently good choice.
@@
We start our discussion by reviewing the notion of acceptance testing for fixed-length protocols. There, before the protocol is executed, one has to determine and fix an expected channel behaviour by determining expected observations for the witness operators. This is usually a single data point. In this case, of course, every tiny deviation from the expected behaviour leads to a deviation from the observed statistics. During the acceptance test, one compares these observed statistics with the expected statistics, and it aborts if they do not match. Therefore, under this ‘unique-acceptance’ scenario, one would abort the protocol almost all the time, making it secure but essentially useless. We can fix this problem by introducing some margin to extend the set of accepted statistics (see parameter $t_{X}$ in Theorem 1). As elaborated on earlier, by doing so, we build a multi-dimensional interval centered on our original expected statistics, which increases the probability of acceptance. This, however, comes at the cost of extending the set of states that pass the acceptance set, hence the set of states that potentially produced the observed statistics, and thus lowers the key rate. To ease the illustration, we parametrize $t_{X}$ for each observable $X\in\Theta$ by a positive real number $t_{F}$, $t_{X}=t_{F}\mu_{X}$. In contrast, the variable-length security argument is built around the observed statistics and not around the expected statistics and adapts the secure key length to the quality of the observations. Therefore, for variable-length key rates, we can set $t_{F}=0$. As derived in Section IV, the acceptance testing security parameter $\epsilon_{\mathrm{AT}}$ differs between the fixed- and the variable-length scenarios, where $\epsilon_{\mathrm{AT}}$ is the sum of the acceptance-testing epsilons $\epsilon_{X}$ for each single observable $X\in\Theta$. To simplify application for the variable-length scenario, from now on, we require that all the $\epsilon_{X}$ are equal, hence $\epsilon_{\mathrm{AT}}=|\Theta|\epsilon_{X}$ and $\mu_{X}=\sqrt{\frac{x^{2}}{2m_{X}}\ln\left(\frac{2}{\epsilon_{X}}\right)}$.
@@
Let us begin by developing an intuition for those two different notions of key rates on fluctuating channels. In Figure 5a, we simulate $100$ runs of the ‘Statistically Fluctuating Channel’ with the following channel parameters $\mu_{\mathrm{sol}}=2\times 10^{4}$, $\sigma_{\mathrm{sol}}=1\times 10^{4}$, $\mu_{\mathrm{loss}}=0.99$ and $\sigma_{\mathrm{loss}}=0.005$. The pair production rate was set to $0.1$ per time frame, the detection efficiency was assumed to be $\eta_{D}=0.90$, and the dark-count rate was assumed to be $100$ per second. Furthermore, we set $t_{F}=0.5$. We plot the secure key rate for each run, for both the fixed-length and for the variable-length protocol. While the fixed-length protocol either produces a key of fixed length or aborts, it can be seen readily how the key lengths for the variable-length protocol adapt to the observed data, while it only fails for two out of $100$ data points, yielding zero key. In the case of acceptance, the difference in key rate between the fixed-length and the variable-length case is (in this example) about $20\%$. However, since for the fixed-length scenario more than half of the runs are not accepted, this results in the average variable-length key rates being roughly $2.5$ times higher than the fixed-length key rates.
@@
In real scenarios, the channel parameters are usually unknown, so, for the present example, we chose $t_{F}$ heuristically. This leaves the question of whether, if the behavior of the channel - including its statistical parameters - were known, one could find the optimal $t_{F}$ for the fixed-length scenario and close the gap in performance. We are going to explore this question in what follows.
@@
In Figure 5b, we examine the key rates for the fixed-length protocol and its acceptance ratio for different values of $t_{F}$. Again, we assumed a ‘Statistically Fluctuating Channel’ with the channel parameters mentioned above and simulated $1000$ runs for each datapoint. The blue dashed line represents the ’one-shot‘ fixed-length key rates, which are the fixed-length key rates upon acceptance, while the light-red dotted curve shows the associated ratio of accepted rounds. The dot-dashed red curve represents the average fixed-length key rate for each data point (which is the product of the one-shot value with the acceptance ratio). For comparison, we also plot the (average) variable-length key rate (black dotted line). Note that the variable-length key rate does not require $t_{F}>0$ and thus remains constant with $t_{F}$. One can see that in this example, the one-shot fixed-length key rates are higher than the variable-length key rates for $t_{F}<0.25$. However, due to the low acceptance ratio, the expected fixed-length key rates always remain significantly lower than the variable-length rates. For higher values of $t_{F}$, the acceptance rate slowly converges toward $1$ while at the same time, the one-shot key rate drops, resulting overall in a decreasing average key rate. The highest fixed-length key rate is observed for $t_{F}=1$. However, the exact location of the maximum depends on the channel parameters. Thus, even in case one manages to determine the optimal $t_{F}$ a priori, the obtained average fixed-length key rates would still be significantly below the variable-length rates.
@@
Already in the case of stable channels with parameters fluctuating statistically around fixed mean values, variable-length key rates outperform their fixed-length equivalents in every aspect. If, in addition, the mean values fluctuate rapidly, the fixed-length acceptance test will only pass in a tiny fraction of all runs, leading to an expected key rate of nearly zero, while the variable-length key adapts to the observed statistics and will produce a nonzero key in many of those cases. We illustrate this in Figure 6, where we present the results of simulations of the the chaotic channel with parameters $(p_{\textrm{loss}},n_{\text{sol}})\in\{0.97,0.98,0.99\}\times\{1\times 10^{4},2\times 10^{4},1\times 10^{5}\}$ (see Appendix C for details), where each channel parameter is chosen randomly according to a uniform distribution and $t_{F}=0.5$. As expected, the fixed-length protocol aborts in the majority of all cases, yielding zero key. In contrast, the variable-length protocol produces a non-negative key for most runs. The average variable-length rates are now about $6.7$ times higher than the average fixed-length key rate. This illustrates that, in real free-space channels, the advantage of the variable-length security argument will be even more striking.
@@ VII Discussion
In recent years, the development of implementations for high-dimensional QKD on the one side and security analyses on the other side has proceeded at a vastly different paces. While highly entangled states have been produced in various degrees of freedom [Kwiat_1997, Barreiro_2005, Martin_2017, Islam_2017, Bulla_2023, Sulimany_2023, Hu_2020, Bouchard_2021, Schneeloch_2019, Ponce_2022, Maxwell_2022], theoretical analyses have lagged lagging behind, either relying on practically infeasible measurements [Doda_2021, Sheridan_2010] or using computationally intensive convex optimization methods [BKPH_2023] that are limited to very low dimensions, despite taking hours to days to compute. While this problem was solved in the asymptotic regime by Ref. [Kanitschar_2024a], the practically relevant finite-size regime for HD-QKD was still missing.
In this work, we close this gap by proving the first composable finite-size security argument for a general high-dimensional QKD protocol against both collective and coherent attacks. The obtained rates converge to the corresponding asymptotic rates [Kanitschar_2024a] in the limit of large $N$. We obtain non-negative key rates already for practically achievable block sizes of $N=10^{8}$ under viable conditions. This shows that HD-QKD is suitable for the loss and noise regimes [Liao_2017, Bedington_2017] present in free-space and LEO satellite QKD links. Additionally, we demonstrate that our method provides key rates comparable to those obtained by state-of-the-art numerical convex optimisation techniques. While the latter are computationally intense and thus limited to the low double-digit dimensions even on clusters [BKPH_2023], our semi-analytic method unlocks genuine high dimensions, providing key rates that surpass existing results in terms of both achievable key rate and maximum tolerable loss.
In addition to common fixed-length key rates, we provide more practically useful variable-length key rates, which build the composability argument around the observed statistics rather than the expected statistics, significantly improving both the practicality and key rates of the protocol. We discuss and compare these two different notions of security to illustrate the advantages of variable-length key rates.
While we used a time-bin implementation for demonstration purposes, our work is general and can be applied to a variety of platforms with little to no adaptation. In summary, this work provides the missing tool for the full security analysis of general high-dimensional QKD protocols and unlocks the practical use of high-dimensionally entangled states in Quantum Key Distribution. The obvious next steps consist of the application to real implementations and further improvements in the key rate.
@@ Acknowledgements.
F.K. thanks Devashish Tupkary for inspiring discussions and precious feedback on an earlier version of this manuscript, Ian George for insightful discussions, and Shlok Nahar and Matej Pivoluska for valuable comments on an earlier version of the manuscript. This work has received funding from the Horizon Europe research and innovation programme under grant agreement No 101070168 (HyperSpace). F.K. gratefully acknowledges support from the Dieberger-Skalicky foundation. The authors acknowledge TU Wien Bibliothek for financial support through its Open Access Funding Programme.
@@ Appendix A Construction of the set $V(\vec{F}^{\mathrm{obs}})$
[…]
@@ Appendix C Simulation of the Atmospheric Channel
In order to demonstrate the advantages of a variable-length security argument, we simulate secure key rates for two different channel behaviors. First, we consider a statistically fluctuating channel, followed by a rapidly fluctuating channel. Both simulations use the noise model from [BKPH_2023] to relate the experimental quantities pair-production rate $c_{\mathrm{Prod}}$, dark-count rate $d_{\mathrm{dark}}$, detection efficiency $\eta_{d}$, channel loss probability $p_{L}$ and number of environmental (solar) photons $n_{\mathrm{sol}}$ to click matrices. For details regarding this noise model, we refer interested readers to Ref. [BKPH_2023].